OpenStudy (samigupta8):

the value of tan20°.tan40°.tan80°

OpenStudy (experimentx):

add tan 60 to make things easier

OpenStudy (samigupta8):

then also wt is the use of it

OpenStudy (experimentx):

woops!! looks like you can't use reflection formula

OpenStudy (anonymous):

by "add" i think @experimentX means "multiply by"

OpenStudy (experimentx):

lol .. yeah sorry.

OpenStudy (anonymous):

just being picky we knew what you meant

OpenStudy (experimentx):

\[ \tan 20 \tan (60 - 20 ) \tan(60 + 20) = \tan 20 \left( \frac{\tan^2 60 - \tan^ 2 20 }{1 - \tan ^2 20 \tan ^2 60 }\right) \] We also know that \[ \tan (3*20) = \tan 60 \] find the value of tan(20) and put is there

OpenStudy (samigupta8):

i got itt....

OpenStudy (experimentx):

there is easier method write it up as. sin(20) sin(40) sin(80) ------------------- cos(20) cos(40) cos(80) then \[\frac{\sin 20 \sin(60 - 20)\sin(60 + 20)}{\cos 20 \cos(60 - 20)\cos(60 + 20)}\] after slight manipulation the numerator should turn to sin(3*20) and denominator should turn into cos(3*20) where you get tan(60) which is your value.1